Hello my name is Stephnanie. I am trying to learn how to take problems in English such as: The output is six less than one-half of the input. and make it into an equation, and make a x,y table. please help thanks
Answer by Flannery(124)

Hi Stephanie, I see you are learning about functions.
Usually the input value is the x value and the output value is the y value.
The phrase you used in the question translates to y=1/2x-6.
That may be written f(x)=1/2x-6 (pronounced function of x is one half of x minus six). It may be expressed as f(x) ONLY if for every input x there is ONLY one output y. Let's make a short table for y=1/2x-6.
when y=0
0=1/2x-6; x=12
when y=1
1=1/2x-6; x=14
when y=2
2=1/2x-6; x=16
when x=1
y=1/2(1)-6; y=-5 1/2
when x=2
y=1/2(2)-6; y=-5
when x=3
y=1/2(3)-6; y=-4 1/2
Those are six sets of values for your equation. They describe a function. If you graphed this set of points on x,y coordinates they would make a straight line with a slope of 1/2 and an x intercept of (0,12) yintercept of (0,-6). I hope that got you started.